OpenStudy (anonymous):

What is the domain of f(x)=x^2-x +5

jimthompson5910 (jim_thompson5910):

This is a polynomial. All polynomials have the same domain: the set of real numbers (since you can plug in any real number for x and get some real number out for y or f(x)) So the domain of f(x) is the set of all real numbers.

OpenStudy (anonymous):

Okay can you explain how to solve this problem because I really have no clue.

jimthompson5910 (jim_thompson5910):

It's already solved and answered

jimthompson5910 (jim_thompson5910):

Now sure what you're looking for exactly

OpenStudy (anonymous):

Okay what if I had a problem that said what is the domain of f(x)=1/-X+3. What would be the answer? Could you explain more.

jimthompson5910 (jim_thompson5910):

You cannot divide by zero So -x+3 cannot be zero This means that x cannot be 3 (since this makes -x+3 equal to 0) So the domain is the set of all real numbers BUT x cannot equal 3.

jimthompson5910 (jim_thompson5910):

I'm assuming the problem is \[\Large f(x) = \frac{1}{-x+3}\]

OpenStudy (anonymous):

Yes thats it.

OpenStudy (anonymous):

Im confused.. why zero?

OpenStudy (anonymous):

If x cannot equal to zero then what is it?

jimthompson5910 (jim_thompson5910):

because you can't divide by zero

jimthompson5910 (jim_thompson5910):

no, x cannot equal 3

jimthompson5910 (jim_thompson5910):

x can be any number but 3

OpenStudy (anonymous):

Ok, if you cannot divide by zero then what do you divide by?

jimthompson5910 (jim_thompson5910):

you can divide by any number but zero

jimthompson5910 (jim_thompson5910):

ex: 2/1 = 2 2/0.5 = 4 but 2/0 is undefined

OpenStudy (anonymous):

Okay I understand that. So 1/-x+3 would be?

OpenStudy (anonymous):

undefined?

jimthompson5910 (jim_thompson5910):

it's only undefined when x = 3

jimthompson5910 (jim_thompson5910):

otherwise, it's some number

OpenStudy (anonymous):

Ok.. well back to the first problem. It asks me to find the domain of f(x)= x^2-x+5. But it doesn't tell me what x is, When I solved it I got this... f(x)= -5 Can you please help me with this problem.

jimthompson5910 (jim_thompson5910):

but the domain of any polynomial is the set of all real numbers

jimthompson5910 (jim_thompson5910):

there has to be more to this

OpenStudy (anonymous):

There isn't. Thats all to the problem.. I have 5 problems like that.

jimthompson5910 (jim_thompson5910):

hmm not sure then

jimthompson5910 (jim_thompson5910):

it sounds like you're either solving for x or evaluating x for a specific value but you're not finding the domain if it's f(x) = -5

jimthompson5910 (jim_thompson5910):

because the domain again is simply "the set of all real numbers"

OpenStudy (anonymous):

Im not sure if thats the answer. i just tried to solve it.. f(x)= x^2-x=5 -5=x^2-x -5=x

jimthompson5910 (jim_thompson5910):

then I would go with "the set of all real numbers" if all they want is the domain

OpenStudy (anonymous):

f(x)= x^2-x+5 -5=x^2-x -5=x

jimthompson5910 (jim_thompson5910):

esp if this a polynomial

OpenStudy (anonymous):

Okay so if I had a problem like f(x)=1/radical x-3 it would be a set of all real numbers?

jimthompson5910 (jim_thompson5910):

no, because we now have a division by a variable expression

jimthompson5910 (jim_thompson5910):

there's a potential to divide by zero

OpenStudy (anonymous):

Okay how would I figure this out? so I divide what by zero?

jimthompson5910 (jim_thompson5910):

The problem is \[\Large f(x) = \frac{1}{\sqrt{x-3}}\] right?

OpenStudy (anonymous):

Yes!

jimthompson5910 (jim_thompson5910):

ok so the denominator cannot be zero

jimthompson5910 (jim_thompson5910):

if it were, then sqrt(x-3) = 0 x - 3 = 0 x = 3 This means that when x = 3, the denominator is zero. This means that you can plug in any other number BUT x = 3 into the function So the domain is the set of all real numbers but x can't equal 3.

OpenStudy (anonymous):

so the answer can be anything but not 3?

jimthompson5910 (jim_thompson5910):

yes that's one way of stating the domain

jimthompson5910 (jim_thompson5910):

it's probably a more compact way to state what numbers are excluded from the domain (which is x = 3)

OpenStudy (anonymous):

she's looking for interval notation when x does not equal zero, you cant have zero in the denominator, so your domain is (-infinity,0] U [0, infinity)